# Analysis of Hierarchical AoII over unreliable channel: A Stochastic Hybrid System Approach

### v1

@@ D-1 Proof of Necessity
II2). If $k_{1}>\lambda$, $k_{2}>\mu$, then we again have (56a), (56b) and (56c). However, the requirement in (56a) cannot be met since
@@
To summarize, we have proved the necessity of the stability conditions consisting of $0<k_{1}<\lambda$, $0<k_{1}<k_{2}<\mu$ and (51a) as given in Theorem 2.
D-2 Proof of Sufficiency
In order to prove the sufficiency of these two conditions, we need to give the ergodicity of the joint state $(\mathbf{q},\mathbf{x})$ also the boundness of the average AoII $x_{b}$. Let $(\mathbf{q}_{0},\mathbf{x}_{0})$ denotes the initial condition at time $t_{0}$ and $\{t_{i}:i\geqslant 0\}$ be the sequence of transition time. On each interval $[t_{i},t_{i+1})$, the AoII at the monitor evolves as
$x_{2}(s)=x_{2}(t_{i})+\int_{t_{i}}^{s}f_{2}^{\prime}(\mathbf{q}(r),\mathbf{x}(r),r)\mathrm{d}r,\ \forall s\in[t_{i},t_{i+1}),$ (60)
according to (3). Taking norms at both sides we have
$||x_{2}(s)||\leqslant||x_{2}(t_{i})||+\int_{t_{i}}^{s}\max\{k_{2}||x_{2}(r)||,c\}\mathrm{d}r,$ (61)
where $c=1$ is the given slope constant at linear growth in our case [^8] It is easy to verify that $c$ can be any positive constant., and we assume that $k_{2}>k_{1}$ without loss of generality. Since $k_{2}||x_{2}(r)||$ is continuous on each interval $[t_{i},t_{i+1})$, one of the following cases must occur:
I1). $k_{2}||x_{2}(r)||<1$. Then we have
$||x_{2}(s)||\leqslant||x_{2}(t_{i})||+t_{i+1}-t_{i}.\forall s\in[t_{i},t_{i+1}),$ (62)
I2). $k_{2}||x_{2}(r)||>1$. Then we have
$||x_{2}(s)||$ $\leqslant||x_{2}(t_{i})||+\int_{t_{i}}^{s}k_{2}||x_{2}(r)||\mathrm{d}r$ (63)
$\leqslant\exp\Big{[}\int_{t_{i}}^{s}k_{2}\mathrm{d}r\Big{]}||x_{2}(t_{i})||,\forall s\in[t_{i},t_{i+1}),$
I3). $k_{2}||x_{2}(r_{0})||=1$ for some $r_{0}\in[t_{i},t_{i+1})$. Then we can leverage the Bellman-Gronwall Lemma on $r\in[r_{0},t_{i+1})$ to obtain
$||x_{2}(s)||$ $\leqslant\exp\Big{[}\int_{r_{0}}^{s}k_{2}\mathrm{d}r\Big{]}||x_{2}(r_{0})||,\ \forall s\in[r_{0},t_{i+1}).$ (64)
Combining I2) and I3), for $\forall s\in[t_{i},t_{i+1})$ we have
$||x_{2}(s)||\leqslant\exp\Big{[}\int_{t_{i}}^{s}k_{2}\mathrm{d}r\Big{]}\max\{||x_{2}(t_{i})||,\frac{1}{k_{2}}\}.$ (65)
$||x_{2}(t_{i+1})||$ $=||\phi^{x}_{l}(\mathbf{\mathbf{q},\mathbf{x},t})||$ (66)
$\leqslant\exp\Big{[}\int_{t_{i}}^{t_{i+1}}k_{2}\mathrm{d}r\Big{]}\max\{||x_{2}(t_{i})||,\frac{1}{k_{2}},0\},$
where $\phi^{x}_{l}$ is the mapping of the continuous state. The inequality follows from Table I since the AoII at the monitor either keeps growing or drops to zero at the transition time $t_{i+1}$.
In the worst case, the second entry of initial condition $\mathbf{x}_{0}=(x_{1}(t_{0}),x_{2}(t_{0}))$ is larger than the constant $\frac{1}{k_{2}}$ and the first term in the maximum function of (64) always dominates as the AoII grows from $i=0$ to $i=k-1$, which yields
$||x_{2}(t_{k})||\leqslant\exp\Big{[}\int_{t_{0}}^{t_{k}}k_{2}\mathrm{d}r\Big{]}||x_{2}(t_{0})||.$ (67)
Denoting $t_{k}$ as the last transition time before some arbitrary time $t>t_{k}$, we obtain
$||x_{2}(t)||$ $\leqslant\exp\Big{[}\int_{t_{k}}^{t}k_{2}\mathrm{d}r\Big{]}\max\{||x_{2}(t_{k})||,\frac{1}{k_{2}}\}$ (68)
$\leqslant\exp\Big{[}\int_{t_{0}}^{t}k_{2}\mathrm{d}r\Big{]}||x_{2}(t_{0})||,\forall t\in[t_{k},t_{k+1}),$
from (64) and (67) for $i=k$ and $s=t$. Note that we have ergodicity of discrete state $\mathbf{q}\in\mathbb{Q}$, which implies the state $(\mathbf{q},x_{2}(t))$ would eventually return to $(0_{c},0)$ at some $t_{i+1}$ in a finite time even as $t\to\infty$. Therefore, $||x_{2}(t)||$ is always bounded in the quadratic hierarchy case, which in turn gives the ergodicity of the joint state $(\mathbf{q},\mathbf{x})$ as well as the boundness of the average AoII $x_{b}$.
Appendix E Derivations of Corollary 3: The Quadratic Hierarchy over Collision Channel
@@
Appendix F Sufficient Stability Conditions of the Quadratic Hierarchy over Collision Channel
The step to prove the ergodicity is the same as the proof 2) in Appendix D and is omitted here. The determinant of $\mathbf{D}$ is obtained in (73) using the definition.
The expected values of AoII at states $1_{m},2_{m}$ are given in (70) as $c_{1}$ and $c_{2}$, both of which are larger than zero. Then we need to calculate the remaining expected values of AoII at states $0_{e},1_{e},1_{u},2_{u}$, which are summarized in (74).
@@ Appendix F Sufficient Stability Conditions of the Quadratic Hierarchy over Collision Channel
[…]
$+\big{[}\mu(m-k_{2})+(\mu+\mu_{c})(\lambda_{c}+\mu_{c}-k_{2})+2\lambda(\mu_{c}-k_{2})-2k_{2}\lambda_{c}\big{]}k^{2}_{1}$
@@
$v^{2}_{0_{e}}$ $=\frac{1}{\Pi}\biggl{\{}c_{2}\mu\mu_{c}A+c_{1}p\lambda\lambda_{c}\mu\mu_{c}\biggr{\}},$ (74)
$v^{2}_{1_{e}}$ $=\frac{1}{\Pi}\biggl{\{}c_{2}\mu(\lambda+\lambda_{c}-k_{1})A+c_{1}p\lambda\lambda_{c}\mu(\lambda+\lambda_{c}-k_{1})\biggr{\}},$
@@
$A$ $=(k_{2}-p\lambda-\mu-\mu_{c})k_{1}-(\lambda_{c}+\mu)k_{2}+(\lambda_{c}+\mu+\mu_{c})\mu+(\lambda_{c}+\mu)p\lambda,$
$B$ $=k_{1}^{2}-(2\lambda+\lambda_{c}+\mu_{c})k_{1}+m\lambda.$
$\Pi^{\prime}$ $=(\lambda+\mu-k_{1})C,$ (75)
$C$ $=2\mu^{2}(\lambda-k_{1})+2k^{2}_{1}\mu-(4\lambda-k_{2})k_{1}\mu-(2\lambda-k_{1})(\lambda-k_{1})k_{2}-k_{2}\lambda\mu$

### v2

@@ E-1 Proof of Necessity
II2). If $k_{1}>\lambda$, $k_{2}>\mu$, we again have (67a), (67b) and (67c). However, the requirement in (67a) cannot be met since
@@
To summarize, we have proved the necessity of the steady-state stability conditions consisting of $0<k_{1}<\lambda$, $0<k_{1}<k_{2}<\mu$ and (62a) as given in Theorem 2.
E-2 Proof of Sufficiency
Following Appendix B.
Appendix F Derivations of Corollary 3
@@
Appendix G Sufficient Stability Conditions of Hierarchy Scheme 3 over the Collision Channel
The determinant of $\mathbf{D}$ is obtained in (75) using the definition. The expected values of AoII at states $1_{m},2_{m}$ are given in (72) as $c_{1}$ and $c_{2}$, both of which are larger than zero. Then we need to calculate the remaining expected values of AoII at states $0_{e},1_{e},1_{u},2_{u}$, which are summarized in (76).
@@ Appendix G Sufficient Stability Conditions of Hierarchy Scheme 3 over the Collision Channel
$\Pi=\biggl\{$ $\lambda\mu^{2}\mu_{c}+(\lambda\mu\mu_{c}-k_{2}\lambda\mu-k_{2}\lambda_{c})(\lambda+\lambda_{c}+\mu_{c})+(k_{2}-\mu-\mu_{c})k_{1}^{3}\ +\big[\mu(m-k_{2})+(\mu+\mu_{c})(\lambda_{c}+\mu_{c}-k_{2})+2\lambda(\mu_{c}-k_{2})-2k_{2}\lambda_{c}\big]k^{2}_{1}$
@@
$v^{2}_{0_{e}}$ $=\frac{1}{\Pi}\biggl\{c_{2}\mu\mu_{c}A+c_{1}p\lambda\lambda_{c}\mu\mu_{c}\biggr\},\ v^{2}_{1_{e}}=\frac{1}{\Pi}\biggl\{c_{2}\mu(\lambda+\lambda_{c}-k_{1})A+c_{1}p\lambda\lambda_{c}\mu(\lambda+\lambda_{c}-k_{1})\biggr\},$ (76)
@@
$A$ $=(k_{2}-p\lambda-\mu-\mu_{c})k_{1}-(\lambda_{c}+\mu)k_{2}+(\lambda_{c}+\mu+\mu_{c})\mu+(\lambda_{c}+\mu)p\lambda,\ B=k_{1}^{2}-(2\lambda+\lambda_{c}+\mu_{c})k_{1}+m\lambda.$
$C=2\mu^{2}(\lambda-k_{1})+2k^{2}_{1}\mu-(4\lambda-k_{2})k_{1}\mu-(2\lambda-k_{1})(\lambda-k_{1})k_{2}-k_{2}\lambda\mu,\ \Pi^{\prime}=(\lambda+\mu-k_{1})C.$ (77)